Dear Math Tutor,
I need help with this problem. I also need this problem by April 23, 2006.
(a-b)raised to the 2 power -3(a-b)-18=0, find the value of a-b-14 if a-b+3=0

Answer by stanbon(48564)

Let a-b=x then your problem becomes:
x^2-3x-18=0
(x-6)(x+3)=0
x=6 or x=-3
Therefore a-b=6 or a-b=0
2nd question:
What is a-b-14 if a-b+3=0
If a-b+2=0
Then a-b==3
Substituting that into a-b-14
you get -3-14=17
So a-b-15=17 if a-b+3=0
Cheers,
Stan H.